method is elimination solve: 3x + 5y = 10
2x + 5y = -5

Answer by ewatrrr(24785)

 
Hi,
Solving by elimination method
3x + 5y = 10
2x + 5y = -5
3x + 5y = 10
-2x - 5y = 5 |subtracting 2nd EQ from the 1st EQ
x = 15 and y = -7 (30+5y = -5, 5y = -35, y = -7)
CHECKING our Answer***
3*15 + 5*(-7) = 45 - 35 = 10